# Uncertainty-Aware Probabilistic Graph Neural Networks for Road-Level Traffic Accident Prediction

### v3

@@ 3 Methodology
In this section, we first introduce the concepts and define the framework for predicting road-level traffic accidents in Section 3.1. We then illustrate the statistical innovations from the TD distribution to ZITD in the context of uncertainty quantification in Section 3.2. Subsequently, we introduce our new STZITD-GNN model, a spatio-temporal probabilistic modelling framework that combines the strengths of GRU and GAT in Section 3.3. The section ends with an in-depth exploration of the loss function employed within our model in Section 3.4. The major notations used in this section are listed in Table 2.
@@
$Y\leavevmode\nobreak\ /\leavevmode\nobreak\ X$ Road-level traffic accident risk score  /  Road-level spatiotemporal traffic feature
@@
$C_{it}$ Number of accidents on road $v_{i}$ at time slot $t$
$l_{it}^{j}$ The $j^{th}$ cardinality values of accident severities in the time slot $t$ for road $v_{i}$
@@ 3.1 Preliminaries and Problem Definition
Definition 1. Road Traffic Accident Risk Score: A traffic accident refers to an accident that involves one or more road users that leads to physical injury, loss of life, or property damage. We denote the set of roads as $V$ of size $N=|V|$, where $v_{i}$ signifies the $i^{th}$ road. The traffic accident risk on the roads is represented as $Y\in\mathbb{R}^{N\times T}$ for a duration of total $T$ periods, and $y_{it}$ denotes the accident risk in the $t^{th}$ time slot for the road $v_{i}$. Thus, $Y_{t}\in\mathbb{N}^{N}$ signifies the accident risk for all roads in the $t^{th}$ time interval, where $y_{it}$ is its constituent element. The calculation of traffic accident risk score $y_{it}$ for the road $v_{i}$ at the time slot $t$ is given by:
@@
where $C_{i,k}^{t}$ denotes the total number of $k$ accidents with the severity level $l_{j}$ at time slot $t$ for road $v_{i}$. We allocate the accident point to its closest road. As per prior research (Wang et al., 2021a, ; Trirat et al.,, 2023), $l$ is assigned the values 1, 2, and 3, representing minor injury, serious injury, and fatal accident severities, respectively. $y_{it}$ equals 0 if there are no traffic accidents.
@@
Problem Definition: Our model aims to predict the future accident risk score in the next $p$ time windows and the confidence interval of the predicted results per road. It is to use historical records $X_{1:t},Y_{1:t}$ and the graph structure $A$ as input data for training, to predict the probabilistic density function $f(Y_{t+1:t+p})$ of the distribution of $Y_{t+1:t+p}$.
@@
Figure 1: The overall framework of STZITD-GNNs. STZITD-GNNs utilise the ST-GNN Encoder $\mathcal{ST}$ (composed of GRU and GAT encoders) to encode the history time window $1:t$ accident risk $Y_{1:t}$ and road features $X_{1:t}$ with the use of the road connection graph in the spatial-temporal embedding of the road $\mathcal{Z}$. After encoding, the four parameter decoders map $\mathcal{Z}$ into the ZITD parameter space and obtain $\pi_{t+1:t+p},\mu_{t+1:t+p},\phi_{t+1:t+p},\rho_{t+1:t+p}$ for the predicted time window $t+1:t+p$, which determine the predicted road traffic accident risk distribution $f(Y_{t+1:t+p})$.
@@ 3.2 From TD Distribution to ZITD Distribution
This subsection provides an in-depth exploration of the progression from the TD to the ZITD distribution in the specific context of road accident prediction. Building on the work of Lord and Mannering, (2010) and Smyth and Jørgensen, (2002), we recognise the effectiveness of TD distribution to address overdispersed data, indicating its potential applicability in multi-step road-level traffic accident prediction.
As illustrated in Section 3.1, when we aggregate accident data (incorporating various severities) from different roads or identical roads over different time intervals, the resulting distribution of accident counts tends to exhibit overdispersion. In detail, the accident value applied to both accident counts and the associated severity - is the prevalence of ’excess’ zeros, implying more zero entries than would be predicted under a corresponding Poisson process. Meanwhile, in nonzero accident values, a clearly left-skewed but long-tailed distribution presents a smaller proportion of higher severity accidents. In these scenarios, the variance of the accident counts outweighs the mean accident count, creating a deviation from the Gaussian distribution, as underlined by the work of Halder et al., (2019) and Wang et al., (2023).
@@ 3.2.1 TD Distribution
In accordance with Equation 1, we designate $y_{k}$ as the accident risk value of the road $i$ on time slot $t$. As a special case of Exponential Dispersion Models (EDMs), the TD model encompasses a broad family of statistical distributions (Tweedie et al.,, 1984; Jørgensen,, 1987; Jiang et al.,, 2023). A random variable that follows the TD distribution has a probability density function $f_{TD}$ (Jørgensen,, 1987), as follows :
@@
In these equations, $\kappa(\theta)^{{}^{\prime}}$ and $\kappa(\theta)^{{}^{\prime\prime}}$ represent the first and second derivatives of $\kappa(\theta)$, respectively. Meanwhile, the function $\mu\geq 0$ works as the mean parameter. Since the mean $E(y_{k})$ and variance $Var(y_{k})$ are dependent, $\kappa(\theta)^{{}^{\prime\prime}}$ can be directly expressed as $V(\mu)$, thereby providing a direct mean-variance relationship (Mallick et al.,, 2022). Additionally, the TD family encompasses a number of vital distributions for distinct index parameters $\rho$ as demonstrated by $V(\mu)=\mu^{\rho}$. These include the Normal ($\rho=0$), Poisson ($\rho=1$), Gamma ($\rho=2$), Inverse Gaussian ($\rho=3$), and Compound Poisson-Gamma distribution ($21<\rho<$ 2) and could be referred to as A. The Compound Poisson-Gamma distribution is particularly applicable, given its focus on parameterizing zero-inflated and long-tailed data, which aligns with the observed characteristics of real traffic accident data.
The frequency of traffic accidents $C_{k}$ adheres to a Poisson distribution represented by $Pois(\lambda)$ with a mean of $\lambda$, expressed as $C_{k}\sim Pois(\lambda)$. This is natural, as frequency often appears as the main prediction target in the previous literature (Zhou et al., 2020b, ; Trirat et al.,, 2023). Currently, the severity of the accident $l_{k}^{(j)}$ is modelled as independent and identically distributed $i.i.d$ Gamma random variables, $C_{k}\perp l_{k}^{(j)}$ and $\forall j$, denoted as $Gamma(\alpha,\gamma)$. These variables have a mean of $\alpha\gamma$ and a variance of $\alpha\gamma^{2}$, such as $l_{k}^{(j)}\stackrel{{\scriptstyle iid}}{{\sim}}\operatorname{Gamma}(\alpha,\gamma)$. $C_{k}$ and $l_{k}^{(j)}$ together form the model of the occurrence of a single accident. Consequently, the traffic accident risk $y_{k}$ is formulated as a Poisson sum of these i.i.d. Gamma random variables. Considering Equation 1 and considering $21<\rho<$ 2, we can finally reformulate the random variable $y_{k}$ as follows:
In the scenario where $C_{k}=0$, we correspondingly have $y_{k}=0$. The probability density at zero for traffic accident risk, therefore, becomes $\mathbb{P}(y_{k}=0)=\mathbb{P}(C_{k}=0)=\exp(-\lambda)$. In contrast, when $C_{k}>0$, $y_{k}$ is deduced from the sum of $C_{k}$ independent Gamma random variables. For clarity, following the work of Zhou et al., 2022a , we re-parameterize the TD distribution, setting $\theta=\mu^{1-\rho}/(1-\rho)$ and $\kappa(\theta)=\mu^{2-\rho}/(2-\rho)$. This results in the revised TD distribution as:
where
In this case, $\mu,\phi$ and $\rho$ are the three parameters that govern the probability and expected value of the risk of traffic accidents. Parameters in the Gamma and Poisson distributions ($\lambda,\alpha,\gamma$) can be obtained as (Kurz,, 2017; Halder et al.,, 2019; Mallick et al.,, 2022), :
@@ 3.2.2 ZITD Distribution
The TD model, introduced in Section 3.2.1, is designed to account for the occurrence of zeros by allowing a positive probability of zero outcomes, as demonstrated by formula $\mathbb{P}(y_{k}=0)=\mathbb{P}(C_{k}=0)=\exp(-\lambda)$. However, this model struggles to accurately represent accident data on roads where zero values overwhelmingly dominate, leading to an extremely unbalanced distribution that can misrepresent the actual accident rate (Mallick et al.,, 2022; Zhou et al., 2022a, ).
To address this issue, we introduce a refinement, known as the Zero-Inflated Tweedie (ZITD) model with Four-parametric compound distributions. The ZITD model includes a distinctive feature, a sparsity parameter ($\pi$) that specifically addresses the skewness of the data characterised by an excess of zeros. The parameter $\pi$ signifies the likelihood of zero inflation, effectively distinguishing between zero and nonzero occurrences. In contrast, the complement probability ($1-\pi$) aligns with the traditional TD, facilitating a balanced representation of the accident risk values across the spectrum.
@@
In the ZITD model, each parameter plays a distinct role in representing road-level accident risks. The sparsity parameter ($\pi$) is adjusted for instances without accidents, significantly influencing the dispersion parameter ($\phi$) to provide a weighted interpretation suitable for left-skewed accident data. Higher $\phi$ values reflect the improved capability of the model to represent scenarios with lower risk accurately. Additionally, the index parameter ($\rho$) is essential to capture the long-tail distribution of accident risk values. Values near 2 for $\rho$ indicate the robust performance of the model in these scenarios.
@@ 3.3 The framework of Spatiotemporal Zero-Inflated Tweedie Graph Neural Networks (STZITD-GNNs)
In relation to the four parameters $\pi,\mu,\phi,\rho$ of the ZITD probability distributions detailed above, we have designed a probabilistic spatio-temporal graph learning module, named STZITD-GNNs. This module extracts the spatiotemporal correlations $\mathcal{Z}=\mathcal{ST}_{\Theta}(X_{1:t},Y_{1:t},A)$ under the assumption of ZITD for each road accident and predicts future results for the subsequent $p$ time windows through $f_{ZITD}(X_{t+1:t+p}|\mathcal{Z})$, as described in Equation 11.
@@
As illustrated in Figure 1, instead of performing two independent estimations, our methodology starts with the deployment of a GRU acting as a temporal encoder. Subsequently, a GAT is implemented for the purpose of spatial encoding. The combined output from the GRU and GAT is utilised to parameterize the ZITD distribution. The proposed approach is shown in Equation 12. The architecture’s encoding component applies a spatio-temporal embedding scheme complemented with an extra sparsity parameter and distribution parameters. The decoding component is responsible for probabilistic estimations of future multi-step accident risk values, which can be construed as the set of parameters of the outcome distribution. In contrast to the principles of variational autoencoders, which restrict latent embedding with a specific Gaussian distribution for the representation of latent variables (Kipf and Welling, 2016b, ; Hamilton,, 2020), we echo the statistical domain by allowing integration of deep learning methodologies into parameter learning.
@@ 3.3.1 Parameter Decoders
The detailed functions of the temporal encoder, GRU and the spatial encoder, GAT are listed in the B and C. In this section, we highlight the parameter decoders in detail. To transmute the acquired spatio-temporal embeddings into ZITD parameter values, we have devised four parameter decoders to operate on $\mathcal{Z}$. Consequently, four parameters, namely $\pi_{t+1:t+p},\mu_{t+1:t+p},\phi_{t+1:t+p},$ and $\rho_{t+1:t+p}$ can be computed from $\mathcal{Z}$, as detailed below:
@@ 3.4.1 ZITD Loss Function
As previously mentioned, our encoding mechanism employs a spatio-temporal embedding architecture complemented by a sparsity parameter. The decoding component involves a probabilistic estimation of future road traffic accidents. To accurately predict these accidents, the overall learning objective can be seen as maximising the log-likelihood function: $\max\log f_{ZITD}(y_{k}|\pi,\mu,\phi,\rho)$. We use the negative likelihood as our loss function to better fit the distribution into the data. Here, $y$ denotes the ground-truth values corresponding to one of the predicted accidents with the parameters $\pi,\mu,\phi,\rho$. The log-likelihood of ZITD is made up of the parts $y=0$ and $y>0$:
@@ 3.4.2 Time analysis & Training algorithms
1:Spatiotemporal road feature $X\in\mathbb{R}^{N\times T\times d}$, road network graph $G=(V,E,A)$, roads traffic accident $Y\in\mathbb{R}^{N\times T}$, model hyperparameters $M,F^{\prime},F$, and total time slot $T$, time window $t$, $p$.
2:Calculate road traffic accident $Y$ via Eq. (1).
@@ 4 Experiments
In this section, we describe the detailed elements of the STZITD-GNNs model by analysing its performance in accident prediction in three boroughs of London, UK, namely Lambeth, Tower Hamlets and Westminster, each characterised by distinct sociodemographic traits, road networks, and traffic accident profiles. The accident data is accessed from the UK STATS19 dataset, maintained by the Department for Transport. It is made up of three interlinked tables that provide information about accidents, their severity, and related vehicle data  [^2] https://www.gov.uk/guidance/road-accident-and-safety-statistics-guidance. We have configured the model to forecast 14 steps ahead, achieving stable temporal horizon results. Our evaluation focusses on two key aspects: first, demonstrating the model’s ability to effectively quantify the uncertainty of road-level traffic accident data, and second, showing its precision in predicting. We emphasise that our model not only forecasts multi-step accident risk in advance but also does so with a higher degree of reliability and precision at the finer granularity compared to several relevant baseline models.
@@
4.1 Traffic Accident Data Description
4.1.1 Statistic Analysis

### v4

@@ 3 Methodology
In this section, we first introduce the concepts and define the framework for predicting road-level traffic crashes in Section 3.1. We then illustrate the statistical innovations from the TD distribution to ZITD in the context of uncertainty quantification in Section 3.2. Subsequently, we introduce our new STZITD-GNN model, a spatio-temporal probabilistic modelling framework that combines the strengths of GRU and GAT in Section 3.3. The section ends with an in-depth exploration of the loss function employed within our model in Section 3.4. The major notations used in this section are listed in Table 2.
@@
$Y\leavevmode\nobreak\ /\leavevmode\nobreak\ X$ Road-level traffic crash risk score  /  Road-level spatiotemporal traffic feature
@@
$C_{it}$ Number of crashes on road $v_{i}$ at time slot $t$
$l_{it}^{j}$ The $j^{th}$ cardinality values of crash severities in the time slot $t$ for road $v_{i}$
@@ 3.1 Preliminaries and Problem Definition
Definition 1. Road Traffic Crash Risk Values: A traffic crash refers to an crash that involves one or more road users that leads to physical injury, loss of life, or property damage. We denote the set of roads as $V$ of size $N=|V|$, where $v_{i}$ signifies the $i^{th}$ road. The traffic crash risk on the roads is represented as $Y\in\mathbb{R}^{N\times T}$ for a duration of total $T$ periods, and $y_{it}$ denotes the crash risk in the $t^{th}$ time slot for the road $v_{i}$. Thus, $Y_{t}\in\mathbb{N}^{N}$ signifies the crash risk for all roads in the $t^{th}$ time interval, where $y_{it}$ is its constituent element. The calculation of traffic crash risk score $y_{it}$ for the road $v_{i}$ at the time slot $t$ is given by:
@@
where $C_{i,k}^{t}$ denotes the total number of $k$ crashes with the severity level $l_{j}$ at time slot $t$ for road $v_{i}$. We allocate the crash point to its closest road. As per prior research (Wang et al., 2021a, ; Trirat et al.,, 2023), $l$ is assigned the values 1, 2, and 3, representing minor injury, serious injury, and fatal crash severities, respectively. $y_{it}$ equals 0 if there are no traffic crashes.
@@
Problem Definition: Our model aims to predict the future crash risk score in the next $p$ time windows and the confidence interval of the predicted results per road. It is to use historical records $X_{1:t},Y_{1:t}$ and the graph structure $A$ as input data for training, to predict the probabilistic density function $f(Y_{t+1:t+p})$ of the distribution of $Y_{t+1:t+p}$.
@@
Figure 1: The overall framework of STZITD-GNNs. STZITD-GNNs utilise the ST-GNN Encoder $\mathcal{ST}$ (composed of GRU and GAT encoders) to encode the history time window $1:t$ crash risk $Y_{1:t}$ and road features $X_{1:t}$ with the use of the road connection graph in the spatial-temporal embedding of the road $\mathcal{Z}$. After encoding, the four parameter decoders map $\mathcal{Z}$ into the ZITD parameter space and obtain $\pi_{t+1:t+p},\mu_{t+1:t+p},\phi_{t+1:t+p},\rho_{t+1:t+p}$ for the predicted time window $t+1:t+p$, which determine the predicted road traffic crash risk distribution $f(Y_{t+1:t+p})$.
@@ 3.2 From TD Distribution to ZITD Distribution
This subsection provides an in-depth exploration of the progression from the TD to the ZITD distribution in the specific context of road crash prediction. Building on the work of Lord and Mannering, (2010) and Smyth and Jørgensen, (2002), we recognise the effectiveness of TD distribution to address overdispersed data, indicating its potential applicability in multi-step road-level traffic crash prediction.
As illustrated in Section 3.1, when we aggregate crash data (incorporating various severities) from different roads or identical roads over different time intervals, the resulting distribution of crash counts tends to exhibit overdispersion. In detail, the crash value applied to both crash counts and the associated severity - is the prevalence of ’excess’ zeros, implying more zero entries than would be predicted under a corresponding Poisson process. Meanwhile, in nonzero crash values, a clearly left-skewed but long-tailed distribution presents a smaller proportion of higher severity crashes. In these scenarios, the variance of the crash counts outweighs the mean crash count, creating a deviation from the Gaussian distribution, as underlined by the work of Halder et al., (2019) and Wang et al., (2023).
@@ 3.2.1 TD Distribution
In accordance with Equation 1, we designate $y_{k}$ as the crash risk value of the road $i$ on time slot $t$. As a special case of Exponential Dispersion Models (EDMs), the TD model encompasses a broad family of statistical distributions (Tweedie et al.,, 1984; Jørgensen,, 1987; Jiang et al.,, 2023). A random variable that follows the TD distribution has a probability density function $f_{TD}$ (Jørgensen,, 1987), as follows :
@@
In these equations, $\kappa(\theta)^{{}^{\prime}}$ and $\kappa(\theta)^{{}^{\prime\prime}}$ represent the first and second derivatives of $\kappa(\theta)$, respectively. Meanwhile, the function $\mu\geq 0$ works as the mean parameter. Since the mean $E(y_{k})$ and variance $Var(y_{k})$ are dependent, $\kappa(\theta)^{{}^{\prime\prime}}$ can be directly expressed as $V(\mu)$, thereby providing a direct mean-variance relationship (Saha et al.,, 2020; Mallick et al.,, 2022). Additionally, the TD family encompasses a number of vital distributions for distinct index parameters $\rho$ as demonstrated by $V(\mu)=\mu^{\rho}$. These include the Normal ($\rho=0$), Poisson ($\rho=1$), Gamma ($\rho=2$), Inverse Gaussian ($\rho=3$), and Compound Poisson-Gamma distribution ($21<\rho<$ 2) and could be referred to as A. The Compound Poisson-Gamma distribution is particularly applicable, given its focus on parameterizing zero-inflated and long-tailed data, which aligns with the observed characteristics of real traffic crash data. The $\mu,\phi$ and $\rho$ are the three parameters that govern the probability and expected value of the risk of traffic crashes, which can be found from B.
@@ 3.2.2 ZITD Distribution
The TD model, introduced in Section 3.2.1, is designed to account for the occurrence of zeros by allowing a positive probability of zero outcomes, as demonstrated by formula $\mathbb{P}(y_{k}=0)=\mathbb{P}(C_{k}=0)=\exp(-\lambda)$. However, this model struggles to accurately represent crash data on roads where zero values overwhelmingly dominate, leading to an extremely unbalanced distribution that can misrepresent the actual crash rate (Mallick et al.,, 2022; Zhou et al., 2022a, ).
To address this issue, we introduce a refinement, known as the Zero-Inflated Tweedie (ZITD) model with Four-parametric compound distributions. The ZITD model includes a distinctive feature, a sparsity parameter ($\pi$) that specifically addresses the skewness of the data characterised by an excess of zeros. The parameter $\pi$ signifies the likelihood of zero inflation, effectively distinguishing between zero and nonzero occurrences. In contrast, the complement probability ($1-\pi$) aligns with the traditional TD, facilitating a balanced representation of the crash risk values across the spectrum.
@@
In the ZITD model, each parameter plays a distinct role in representing road-level crash risks. The sparsity parameter ($\pi$) is adjusted for instances without crashes, significantly influencing the dispersion parameter ($\phi$) to provide a weighted interpretation suitable for left-skewed crash data. Higher $\phi$ values reflect the improved capability of the model to represent scenarios with lower risk accurately. Additionally, the index parameter ($\rho$) is essential to capture the long-tail distribution of crash risk values. Values near 2 for $\rho$ indicate the robust performance of the model in these scenarios.
@@ 3.3 The framework of Spatiotemporal Zero-Inflated Tweedie Graph Neural Networks (STZITD-GNNs)
In relation to the four parameters $\pi,\mu,\phi,\rho$ of the ZITD probability distributions detailed above, we have designed a probabilistic spatio-temporal graph learning module, named STZITD-GNNs. This module extracts the spatiotemporal correlations $\mathcal{Z}=\mathcal{ST}_{\Theta}(X_{1:t},Y_{1:t},A)$ under the assumption of ZITD for each road crash and predicts future results for the subsequent $p$ time windows through $f_{ZITD}(X_{t+1:t+p}|\mathcal{Z})$, as described in Equation 7.
@@
As illustrated in Figure 1, instead of performing two independent estimations, our methodology starts with the deployment of a GRU acting as a temporal encoder. Subsequently, a GAT is implemented for the purpose of spatial encoding. The combined output from the GRU and GAT is utilised to parameterize the ZITD distribution. The proposed approach is shown in Equation 8. The architecture’s encoding component applies a spatio-temporal embedding scheme complemented with an extra sparsity parameter and distribution parameters. The decoding component is responsible for probabilistic estimations of future multi-step crash risk values, which can be construed as the set of parameters of the outcome distribution. In contrast to the principles of variational autoencoders, which restrict latent embedding with a specific Gaussian distribution for the representation of latent variables (Kipf and Welling, 2016b, ; Hamilton,, 2020), we echo the statistical domain by allowing integration of deep learning methodologies into parameter learning.
@@ 3.3.1 Parameter Decoders
The detailed functions of the temporal encoder, GRU and the spatial encoder, GAT are listed in the C and D. In this section, we highlight the parameter decoders in detail. To transmute the acquired spatio-temporal embeddings into ZITD parameter values, we have devised four parameter decoders to operate on $\mathcal{Z}$. Consequently, four parameters, namely $\pi_{t+1:t+p},\mu_{t+1:t+p},\phi_{t+1:t+p},$ and $\rho_{t+1:t+p}$ can be computed from $\mathcal{Z}$, as detailed below:
@@ 3.4.1 ZITD Loss Function
As previously mentioned, our encoding mechanism employs a spatio-temporal embedding architecture complemented by a sparsity parameter. The decoding component involves a probabilistic estimation of future road traffic crashes. To accurately predict these crashes, the overall learning objective can be seen as maximising the log-likelihood function: $\max\log f_{ZITD}(y_{k}|\pi,\mu,\phi,\rho)$. We use the negative likelihood as our loss function to better fit the distribution into the data. Here, $y$ denotes the ground-truth values corresponding to one of the predicted crashes with the parameters $\pi,\mu,\phi,\rho$. The log-likelihood of ZITD is made up of the parts $y=0$ and $y>0$:
@@ 3.4.2 Time analysis & Training algorithms
1:Spatiotemporal road feature $X\in\mathbb{R}^{N\times T\times d}$, road network graph $G=(V,E,A)$, roads traffic crash $Y\in\mathbb{R}^{N\times T}$, model hyperparameters $M,F^{\prime},F$, and total time slot $T$, time window $t$, $p$.
2:Calculate road traffic crash $Y$ via Eq. (1).
@@ 4 Experiments
In this section, we describe the detailed elements of the STZITD-GNNs model by analysing its performance in crash prediction in three boroughs of London, UK, namely Lambeth, Tower Hamlets and Westminster, each characterised by distinct sociodemographic traits, road networks, and traffic crash profiles. The crash data is accessed from the UK STATS19 dataset, maintained by the Department for Transport. It is made up of three interlinked tables that provide information about crashes, their severity, and related vehicle data  [^2] https://www.gov.uk/guidance/road-crash-and-safety-statistics-guidance. We have configured the model to forecast 14 steps ahead, achieving stable temporal horizon results. Our evaluation focusses on two key aspects: first, demonstrating the model’s ability to effectively quantify the uncertainty of road-level traffic crash data, and second, showing its precision in predicting. We emphasise that our model not only forecasts multi-step crash risk in advance but also does so with a higher degree of reliability and precision at the finer granularity compared to several relevant baseline models.
@@
4.1 Traffic Crash Data Description